OpenStudy (anonymous):

Trig Identities: 1+cosθ/1-cosθ=1+1/cosθ/1/cosθ-1

OpenStudy (anonymous):

thats not an identity are you trying to solve ?

OpenStudy (anonymous):

that identity can be solved but you cant find an identity for it

OpenStudy (anonymous):

huh? its like prove each identity

OpenStudy (anonymous):

then that equation

OpenStudy (anonymous):

Can you rewrite that using the equation button?

OpenStudy (anonymous):

if its just solving then answer 0=0+2 (pi sign) n, (pi sign) +2 (pi sign) n

OpenStudy (anonymous):

\[1+\cos \Theta \div 1-\cos \theta =1+(1\div1\div \cos \theta)\div(1\div1 \cos \theta)-1\]

OpenStudy (anonymous):